OpenStudy (anonymous):

Darnell wants to buy ready-mix concrete for his patio. The concrete will fill a space that is 54 feet by 16 feet by 6 inches. A. How many cubic yards of ready-mix cibcrete will Darnell need? 1 cubic yard =27 cubic feet B. Ready-mix concrete cosrs $75.00 per cubic yard. How much will the ready-mix concrete cost for Darnell's patio?

OpenStudy (anonymous):

@xok

OpenStudy (anonymous):

lool

OpenStudy (anonymous):

@xokateoxo

OpenStudy (anonymous):

@xokatexo

OpenStudy (accessdenied):

The space would look like this (not really any scale):  You could start by converting 6 inches into feet. \[ \frac{ 6~\text{inches} }{1} \times \frac{1~\text{foot}}{12~\text{inches}}\\~\\ = \frac{1}{2} \text{feet} \] Then multiply all the dimensions together to get the volume in cubic feet